OpenStudy (anonymous):

Can someone please check my answers .Thanks! 1.)convert the angle theta =40deg 10minutes 25 seconds to a decimal in degrees.round to 2 decimal places i got 40.173--->40.28 2)convert 18.255 degrees to DMS form and then round to the nearest second. I got 18 deg 15minutes 18 second 3.)given theta =53 degreees find one positive and one negative coterminal of theta. I got 307 and -307

OpenStudy (anonymous):

They Look Pretty Legit to Me /.\ Then Agian You Might Wanna Ask @FutureMathProfessor

OpenStudy (anonymous):

k thanks.

OpenStudy (anonymous):

Your Very Welcome (:

OpenStudy (anonymous):

@FutureMathProfessor can you check my work please?

OpenStudy (anonymous):

You are a beast at mathematics. Keep up the good work!

OpenStudy (anonymous):

#1 I got 40.1736; rounding to 2 decimals = 40.17 (.28 is one key to the right. Typo?) #2 We agree on this one #3 is wrong: Your neg coterminal is 53-360 = -307 (this one you got okay) Your pos coterminal is 53+360 = 413 (yes, really. :-) http://hotmath.com/hotmath_help/topics/coterminal-angles.html